# Representing local protein environments with atomistic foundation models

### v2

@@ F.4 Computation details
Appendix G Using likelihoods to define environment similarity measure
G.1 Similarity metric
We analyzed a large collection of protein chemical environments by comparing them using MACE embeddings. For each attribute—such as amino acid type or secondary structure—we estimated a likelihood around each embedding via kernel density estimation with a fixed bandwidth as described in Section 5. These likelihoods were converted into distances using a scaled negative exponential, where higher similarity corresponded to smaller distances. Since each pairwise distance arises from a distribution—based on multiple environments belonging to the same attribute—we applied stochastic multidimensional scaling (MDS) [56, 8], which takes into account both the mean and standard deviation of these distances. The resulting embeddings revealed meaningful clustering that not only separated different amino acid types and secondary structures but also captured finer relationships within them. Chemically or structurally related amino acids and secondary structures exhibited higher similarity, demonstrating that the MACE-based comparison effectively reflects the underlying chemical and structural properties. Additional details and visualizations are provided in Figure A.12.
Appendix H Details of the chemical shift prediction experiments
@@
H.1 Model architecture
We used the same architecture as before (Appendix F.3) with slightly different hyperparameters as described below.
@@ H.2 Computation details
Appendix I Extended AFM interpretations
I.1 Inverting MACE embeddings
In a similar spirit to Maddipatla et al. [44], we investigate whether MACE features can serve not only as representations of protein structures but also as tools to decode them. Specifically, we are interested in using the information contained in the descriptors to select between alternate conformers (altlocs) within a single protein. We therefore treat the MACE features as latent representations from which protein conformers can be reconstructed. As a test case, we examine the PDB structure 4OLE:B resolved to $2.52$, which contains a region of $9$ amino acids ($60-68$) modeled as a superposition of two alternate conformations – one helix and the other a loop. Starting from the helix, we iteratively optimize the positions to align the MACE features of the loop conformation, thereby assessing the potential of these descriptors for structure recovery and conformational transitions.
@@ I.1 Inverting MACE embeddings
The optimization loss function quantifies the discrepancy between the descriptors of the current and target structures and is computed for each atom within the optimized region. In particular, we define each local atomic environment as all atoms within a $5$ radius around the atom of interest. The resulting descriptors capture both geometric and chemical context at multiple equivariant orders. The loss is composed of three terms, each corresponding to a different equivariant order:
$\mathcal{L}=$ $\alpha\,\left\|\mathbf{B}_{i,\eta_{\nu}k,00}-\mathbf{B}^{\text{target}}_{i,\eta_{\nu}k,00}\right\|_{2}+\beta\,\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{B}_{i,\eta_{\nu}k,1M}^{i}-\mathbf{B}_{i,\eta_{\nu}k,1M}^{\text{target},i}\right\|_{2}$
$+\gamma\sum_{i}\left\|\mathbf{B}_{i,\eta_{\nu}k,2M}^{i}-\mathbf{B}_{i,\eta_{\nu}k,2M}^{\text{target},i}\right\|_{F}$
@@
where $\mathbf{B}_{i,\eta_{\nu}k,LM}^{i}$ and $\mathbf{B}_{i,\eta_{\nu}k,LM}^{\text{target}}$ denote the $L$-th order MACE embeddings of the optimized and target conformations, computed as described in Equation 3. For a given $L$, the Clebsch-Gordan coefficients $C_{\eta_{\nu},\mathbf{lm}}^{LM}$ select the valid combinations of input angular momenta. Note that while the second and the third terms are not rotation-invariant, global alignment of the source configuration with the target ensures validity of the comparison.
This combined loss (weighted by $\alpha,\beta,\gamma$) drives the structural optimization by progressively aligning the descriptor representations of the initial and target altloc conformations.
@@
Fig. A.13 illustrates the result of this optimization. While the backbone atom positions are recovered with high accuracy, the side chains of glutamic acid (GLU) $62$ and lysine (LYS) $63$ adopt incorrect orientations. We hypothesize that this discrepancy arises because MACE features primarily capture the local chemical environment within a given radius. Within this range, distinct side-chain rotamers can produce similar local environments, resulting in ambiguous representations. As a result, the descriptors lack the resolution necessary to distinguish between different side-chain conformations.
@@
These results suggest that while MACE features reliably capture the backbone geometry, are less sensitive to side-chain orientations, as distinct rotameric states can yield similar local environments. Recovering both backbone and side-chain configurations would require further exploration. Success in this direction could enable the use of MACE features (and other atomistic features) for coarse-grained protein modeling with minimal loss of structural fidelity.

### v3

@@ H.1 Computation details
H.2 Similarity metric
We analyzed a large collection of protein chemical environments by comparing them using MACE embeddings. For each attribute – such as amino acid type or secondary structure – we estimated a likelihood around each embedding via kernel density estimation with a fixed bandwidth as described in Section 5. These likelihoods were converted into distances using a scaled negative exponential, where higher similarity corresponded to smaller distances. Since each pairwise distance arises from a distribution – based on multiple environments belonging to the same attribute – we applied stochastic multidimensional scaling (MDS) (Rosenberg et al., 2022; Boyarski & Bronstein, 2021), which takes into account both the mean and standard deviation of these distances. The resulting embeddings revealed meaningful clustering that not only separated different amino acid types and secondary structures but also captured finer relationships within them. Chemically or structurally related amino acids and secondary structures exhibited higher similarity, demonstrating that the MACE-based comparison effectively reflects the underlying chemical and structural properties. Additional details and visualizations are provided in Figure A.13.
Appendix I Chemical shift prediction
I.1 Extended physics-grounded chemical shift results
In Section 6 we presented one representative case study to examine whether the shift predictor correctly captures local structural influences on chemical shifts. Here, we provide two additional case studies to further validate these findings. Both probe distinct types of conformational variation and demonstrate that the proposed predictor consistently produces physically meaningful chemical shift trends.
Alternative conformations. We evaluate our shift predictor over a 100 ns MD simulation of the protein 4OLE:B from Rosenberg et al. (2024), which features two stable conformers: conformer A containing a helix and conformer B with a linearly structured loop (Fig. A.12). Our method captures distinct secondary chemical shift distributions for lysine $63$ CA and CB atoms. The helical conformer exhibits the expected increase in CA secondary shift and decrease in CB secondary shift.
I.2 Details of the chemical shift prediction experiments
@@
I.3 Model architecture
We used the same architecture as before (Appendix G.3) with slightly different hyperparameters as described below.
@@ I.4 Computation details
Appendix J Extended MLFF interpretations
J.1 Helix unfolding
As a complement to the main-text case study in Section 7, we analyze the behavior of MLFF embeddings during a helix-to-strand transition. Using MD simulations of a 7-residue peptide (Fig. A.9), we perform PCA on the MACE embeddings of CA atoms in residues $6$ and $7$, and observe that they follow smooth, non-intersecting trajectories. These trajectories reflect the gradual unfolding process, indicating that the descriptor space encodes continuous and interpretable structural transitions.
[…]
J.2 Inverting MACE embeddings
In this section, we explore whether MACE descriptors can be inverted to recover protein structures and distinguish between alternate conformations (altlocs). We treat MACE features not only as structural representations but also as potential tools to decode conformational states. We present two complementary approaches: (i) direct optimization of atom positions to fit MACE embeddings, and (ii) in line with  Maddipatla et al. (2025), guiding AlphaFold3’s reverse diffusion process using MACE descriptors.
J.2.1 Direct optimization of atomic coordinates
We first investigate whether MACE features can serve as latent representations from which protein conformers can be reconstructed. As a test case for descriptor-based structure inversion, we examined the protein 4OLE:B, which contains a nine–residue segment (residues 60606868) modeled as two distinct alternate conformations. In addition to this example, we applied the same analysis to the proteins listed in Table B.15: 3V3S:B, 7EC8:A, 5G51:A, and 4NPU:B. Each of these structures contains a short region annotated with alternate conformations, typically involving subtle backbone displacements rather than large-scale structural rearrangements. For every protein, we optimized the atomic coordinates in the altloc region so that the local MACE descriptors of one conformation match those of the other, enabling us to assess how well the descriptors capture and recover experimentally observed conformational differences. The resulting optimized structures are shown in Fig. A.16 and Fig. A.15.
@@
The optimization loss function quantifies the discrepancy between the descriptors of the current and target structures and is computed for each atom within the optimized region. In particular, we define each local atomic environment as all atoms within a $5$ Å radius around the atom of interest. The resulting descriptors capture both geometric and chemical context at multiple equivariant orders. The loss is composed of three terms:
[…]
$+\gamma\sum_{i}\left\|\mathbf{B}_{i,\eta_{\nu}k,2M}^{i}-\mathbf{B}_{i,\eta_{\nu}k,2M}^{\text{target},i}\right\|_{F},$
@@
where $\mathbf{B}_{i,\eta_{\nu}k,LM}^{i}$ and $\mathbf{B}_{i,\eta_{\nu}k,LM}^{\text{target}}$ denote the $L$-th order MACE embeddings of the optimized and target conformations, computed as described in Equation 3. For a given $L$, the Clebsch-Gordan coefficients $C_{\eta_{\nu},\mathbf{lm}}^{LM}$ select the valid combinations of input angular momenta. While the second and third terms are not rotation-invariant, global alignment of the source configuration with the target ensures validity of the comparison.
@@
This combined loss (weighted by $\alpha,\beta,\gamma$) drives the structural optimization by progressively aligning the descriptor representations of the initial and target altloc conformations. As illustrated in Fig. A.16, the backbone atom positions are recovered with high accuracy, but the side chains of glutamic acid (GLU 62) and lysine (LYS 63) adopt incorrect orientations. We hypothesize that this discrepancy arises because MACE features primarily capture the local chemical environment within a given radius, which cannot always resolve distinct side-chain rotamers.
@@
These results suggest that while MACE features reliably capture backbone geometry, they are less sensitive to side-chain orientations. Recovering both backbone and side-chain configurations would require further exploration, but success in this direction could enable the use of MACE features (and other atomistic embeddings) for coarse-grained protein modeling with minimal loss of structural fidelity.
J.2.2 Guiding AlphaFold3 reverse diffusion
We next investigate whether MACE descriptors can guide AlphaFold3’s diffusion-based structure generation process. Specifically, we focus on proteins with alternate conformations (altlocs)—structural heterogeneity intrinsic to the protein within the unit cell. We analyze a subset of four proteins (Table B.15) exhibiting distinct backbone-separated altlocs, from Rosenberg et al. Rosenberg et al. (2024).
Starting from Gaussian noise, we guide AlphaFold3’s reverse diffusion process to recover a target protein structure that minimizes a descriptor-based loss $\mathcal{L}$. The guided reverse SDE Song & Ermon (2019) is given by
$d\mathbf{X}=-\Big(\tfrac{1}{2}\mathbf{X}+\nabla_{\mathbf{X}}\log p_{t}(\mathbf{X}\mid\mathbf{a})+\eta\nabla_{\mathbf{X}}\mathcal{L}\Big)\beta_{t}dt+\sqrt{\beta_{t}}\mathbf{N},$
where $\mathbf{X}$ are atomic coordinates, $\mathbf{a}$ is the amino acid sequence, $\beta_{t}$ is the noise schedule, and $\mathbf{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The term $\nabla_{\mathbf{X}}\log p_{t}(\mathbf{X}\mid\mathbf{a})$ is the unconditional score function, while $\nabla_{\mathbf{X}}\mathcal{L}$ provides descriptor-based guidance. The hyperparameter $\eta$ scales the guidance term.
To invert MACE embeddings, we define $\mathcal{L}$ as a discrepancy between the descriptors of the diffusion variable $\mathbf{X}$ and those of the target conformation $\mathbf{X}^{\text{target}}$ in the altloc region. The atomic environment is specified as all atoms within a $5$ Å radius:
$\mathcal{L}=-\begin{cases}\dfrac{1}{2}\left\|\mathbf{B}_{0}(\mathbf{X}_{t})-\mathbf{B}_{0}^{\text{target}}(\mathbf{X^{\mathrm{target}}})\right\|_{2}^{2},&\text{if }\left\|\mathbf{B}_{0}(\mathbf{X}_{t})-\mathbf{B}_{0}^{\text{target}}(\mathbf{X^{\mathrm{target}}})\right\|_{1}\leq\delta,\\[12.0pt] \delta\left(\left\|\mathbf{B}_{0}(\mathbf{X}_{t})-\mathbf{B}_{0}^{\text{target}}(\mathbf{X^{\mathrm{target}}})\right\|_{1}-\tfrac{1}{2}\delta\right),&\text{otherwise.}\end{cases}$ (14)
Here, $\mathbf{B}_{0}$ is the zeroth-order invariant descriptor derived from $\mathbf{X}_{t}$, and $\mathbf{B}_{0}^{\text{target}}$ is the corresponding descriptor of the target conformation. The threshold $\delta$ switches the loss from L2 (small errors) to L1 (large errors). For all structures, we set $\eta=0.5$ and $\delta=0.25$.
As shown in Fig. A.14 A, unguided AlphaFold3 typically captures only one of the two alternate conformers. By guiding the reverse diffusion process with MACE descriptors, we can recover the alternate conformation that AlphaFold3 misses or preserve its original prediction. To quantitatively assess the distribution, we compute the normalized distance of each generated structure $\mathbf{X}$ relative to conformations $\mathbf{X}^{\mathrm{A}}$ and $\mathbf{X}^{\mathrm{B}}$ modeled in the PDB (Burley et al., 2017):
$d_{\text{A}}$ $=\left\|\mathbf{X}-\mathbf{X}^{\text{A}}\right\|_{2}^{2}$
$d_{\text{B}}$ $=\left\|\mathbf{X}-\mathbf{X}^{\text{B}}\right\|_{2}^{2}$
Normalized Distance $=\left[1-\min\left(\frac{d_{\text{B}}}{d_{\text{A}}},\frac{d_{\text{A}}}{d_{\text{B}}}\right)\right]\cdot\text{sign}(d_{\text{A}}-d_{\text{B}}).$ (15)
Negative values indicate proximity to conformation A, while positive values indicate proximity to conformation B. The resulting distributions (Fig. A.14 B) again show that guiding the diffusion process with descriptors shifts the distribution toward the mode AlphaFold3 misses or preserves the original distribution. Both results indicate that MACE descriptor-guided diffusion reliably steers AlphaFold3 toward accurate backbone geometries of both conformations in most cases.
Appendix K Use of Large Language Models (LLMs)
The use of LLMs was restricted to minor proof-reading, stylistic polishing, and correcting typos in the text. All intellectual contributions, experiments, data analyses, and conclusions were executed and verified exclusively by the authors. Responsibility for the scientific content rests with the authors, and no part of the research process was delegated to LLMs.